### v1

@@ 4 Numerical Results
In this section, we compare the numerical performance of the proposed GRTR and LMNegCur algorithms with the gradient descent (GDA) algorithm, the MCN algorithm Luo2022FindingSS , the MINIMAX-TR algorithm, and the MINIMAX-TRACE algorithm Yao2024TwoTR in solving a synthetic minimax optimization problem and an adversarial deep learning problem. The first experiment is implemented using Python 3.6 on a laptop with an Intel Core i7 2.8GHz and 8GB of RAM, while the second experiment is conducted on a 2.2GHz CPU, 80GB of RAM, and an Nvidia RTX 4090D GPU.
4.1 A synthetic minimax optimization problem
We consider the following nonconvex-strongly concave minimax problem Yao2024TwoTR :
@@ 4.1 A synthetic minimax optimization problem
In the experiment, we set the number of inner iterations $N_{t}$ to $1000$, the dimension of the variable $y$ to $5$, and the dimension of the variable $x$ to $\{10,20\}$. The initial point $x_{0}$ is set to $(10^{-3},\cdots,10^{-3})^{\top}$, close to the stationary point $(0,\cdots,0)^{\top}$, and $y_{0}$ is a random vector. For the problem (4.1), the parameter $L$ is $\{1,1.5,2\}$, and the parameter $\gamma$ is $1$. Figure 1 shows the numerical performance of the six tested algorithms for solving the problem (4.1) under different parameter selections, where the horizontal axis represents the running time and the vertical axis represents the difference between $P(x)$ and the optimal value $P^{*}$.
Since there are $n$ saddle points in problem (4.1), and it is difficult for the GDA algorithm to escape from the saddle point when it is trapped in the saddle point, this also causes the curve of the GDA algorithm to present multiple horizontal line segments. As can be seen from Fig. 1, all second-order algorithms can effectively escape from the saddle point. The MINIMAX-TR algorithm converges slowly because the trust region radius is set to a small value. The proposed GRTR algorithm and LMNegCur algorithm are both better than the state-of-the-art trust region algorithm MINIMAX-TRACE and the cubic regularization algorithm MCN. Numerical results show the effectiveness of the proposed GRTR algorithm and LMNegCur algorithm.
Figure 1: Numerical results of the six tested algorithms for solving (4.1) with different choices of $n$, $L$ and $\gamma$.
@@ 4.2 An adversarial deep learning problem
We compare the proposed GRTR and LMNegCur algorithms with the GDA, MINIMAX-TR, and MCN algorithms. For all algorithms, we set the batch size to $64$. We select the gradient Lipschitz constant $L_{1}$ from $\{1,5,10\}$ and the Hessian Lipschitz constant $L_{2}$ from $\{10,50,100\}$ for experiments. The experimental results showing the most stable and optimal performance are then shown. For the GRTR algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $\sigma=\sqrt{10}$, $r=1/\sqrt{10}$, and $\epsilon=0.01$, and solve the trust region subproblem by iterating $10$ times Curtis2021TrustN the truncated conjugate gradient (CG) method. For the LMNegCur algorithm and the MINIMAX-TR algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $L_{2}=50$, and $\epsilon=0.01$. For the GDA algorithm, we set the step size of both $x$ and $y$ to 0.01. In addition, for the MCN algorithm, we solve the subproblem about ${y}$ by $30$ gradient descent iterations with a step size of $0.01$, which is the same as the GRTR algorithm and the LMNegCur algorithm. In addition, the cubic subproblem is solved by Algorithm 4 in Luo2022FindingSS , where $L=2$, $M=50$, $\sigma=0.001$, and $\mathcal{K}(\epsilon,\delta^{{}^{\prime}})=10$.
Figure 2 shows the numerical performance of the five tested algorithms when solving problem (4.2), where in sub-figure (a), the horizontal axis represents the running time and the vertical axis represents the test accuracy; while in sub-figure (b), the horizontal axis represents the number of iterations $t$ and the vertical axis represents the function value $P(x)$. It can be seen that the number of iterations and convergence time of all second-order algorithms are significantly less than that of the GDA algorithm, while GRTR performs slightly better than other second-order algorithms. These results show the advantages of the two proposed algorithm.
@@ 5 Conclusions
In this paper, we propose a gradient norm regularized trust region (GRTR) algorithm and a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur). Both the GRTR algorithm and the LMNegCur algorithm can achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$. Extending the proposed algorithms to achieve optimal iteration complexity but with inexact subproblem solutions deserves further investigation.

### v2

@@ 4 Numerical Results
In this section, we compare the numerical performance of the proposed GRTR (IGRTR) and LMNegCur (ILMNegCur) algorithms with the gradient descent (GDA) algorithm, the MCN (IMCN) algorithm Luo2022FindingSS , the MINIMAX-TR algorithm, and the MINIMAX-TRACE algorithm Yao2024TwoTR in solving a synthetic minimax optimization problem (a highly artificial piecewise function with smooth components connecting the segments) and a nonconvex-strongly concave minimax problem with sinusoidal perturbation and an adversarial deep learning problem. The first two experiments are implemented using Python 3.6 on a laptop with an Intel Core i7 2.8GHz and 8GB of RAM, while the last experiment is conducted on a 2.2GHz CPU, 80GB of RAM, and an Nvidia RTX 4090D GPU.
4.1 Synthetic minimax optimization problem
We first consider the following nonconvex-strongly concave minimax problem Yao2024TwoTR :
@@ 4.1 Synthetic minimax optimization problem
In the experiment, we set the number of inner iterations $N_{t}$ to $1000$, the dimension of the variable $y$ to $5$, and the dimension of the variable $x$ to $\{10,20\}$. The initial point $x_{0}$ is set to $(10^{-3},\cdots,10^{-3})^{\top}$, close to the stationary point $(0,\cdots,0)^{\top}$, and $y_{0}$ is a random vector. For the problem (4.1), the parameter $L$ is $\{1,1.5,2\}$, and the parameter $\gamma$ is $1$. Figure 1 shows the numerical performance of the nine tested algorithms for solving the problem (4.1) under different parameter selections, where the horizontal axis represents the running time and the vertical axis represents the difference between $P(x)$ and the optimal value $P^{*}$.
Since there are $n$ saddle points in problem (4.1), and it is difficult for the GDA algorithm to escape from the saddle point when it is trapped in the saddle point, this also causes the curve of the GDA algorithm to present multiple horizontal line segments. As can be seen from Fig. 1, all second-order algorithms can effectively escape from the saddle point. The MINIMAX-TR algorithm converges slowly because the trust-region radius is set to a small value. The proposed GRTR algorithm and LMNegCur algorithm are both better than the state-of-the-art trust-region algorithm MINIMAX-TRACE and the cubic regularization algorithm MCN. The ILMNegCur algorithm performs better than other second-order algorithms because the approximate minimum eigenvalue of the Hessian matrix is achieved very early in the iterations of the Lanczos process, while the IGRTR and IMCN algorithms require more iterations and running time to escape from the saddle point. Numerical results show the effectiveness of the proposed GRTR (IGRTR) algorithm and LMNegCur (ILMNegCur) algorithm.
Figure 1: Numerical results of the tested algorithms for solving (4.1) with different choices of $n$, $L$ and $\gamma$.
4.2 Nonconvex-strongly concave problem with sinusoidal perturbation
We consider the following nonconvex-strongly concave problem with sinusoidal perturbation Zhang2024AGDA :
$f(x,y)=\sin(\sqrt{L-1}\cdot\sqrt{\|x\|^{2}+1})+\frac{1}{2}x^{\top}Qx+x^{\top}Ay-\frac{\mu_{y}}{2}\|y\|^{2}$ (4.2)
where $A\in\mathbb{R}^{n\times m}$, $Q\in\mathbb{R}^{n\times n}$, and $\mu_{y}>0$.
In the experiment, we set $m=n\in\{100,1000\}$, $\mu_{y}=1$ and $L\in\{5,20\}$, $A=V\Lambda_{A}V^{-1}$ and $Q=V\Lambda_{Q}V^{-1}$, where $V$ is an orthogonal matrix, $\Lambda_{Q}=\frac{\Lambda_{Q}^{0}}{\|\Lambda_{Q}^{0}\|_{2}}$, where $\Lambda_{Q}^{0}$ is a random diagonal matrix whose diagonal elements are uniformly randomly sampled from the interval $[-1,1]$, and $\Lambda_{A}$ satisfies $\Lambda_{Q}+\frac{1}{\mu_{y}^{2}}\Lambda_{A}^{2}\succeq 0$. It is not difficult to verify that $f(\cdot,\cdot)$ is a $L$-gradient Lipschitz continuous function. Figure 2 shows the numerical performance of nine test algorithms for solving the problem (4.2) under different parameter selections, where the horizontal axis represents the running time and the vertical axis represents the gradient norm of $P(x):=\max_{y}f(x,y)$. We run each algorithm until $\|\nabla P(x_{t})\|\leqslant 1e-5$.
The results show that the GDA algorithm still converges relatively slowly. Although the second-order methods have a higher cost per iteration, they still outperform the GDA algorithm in terms of running time, especially in converging to a high-quality solution.
Figure 2: Numerical results of the nine tested algorithms for solving (4.2) with different choices of $n$ and $L$.
@@ 4.3 An adversarial deep learning problem
Since exact computation of the second-order subproblems can be a computationally expensive step for large-scale problems, we only compare the proposed IGRTR and ILMNegCur algorithms with the GDA, IMCN, and IMINIMAX-TR algorithms, which are inexact versions of the algorithms. For all algorithms, we set the batch size to $64$. We select the gradient Lipschitz constant $L_{1}$ from $\{1,5,10\}$ and the Hessian Lipschitz constant $L_{2}$ from $\{10,50,100\}$ for experiments. The experimental results showing the most stable and optimal performance are then shown. For the IGRTR algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $\sigma=\sqrt{10}$, $r=1/\sqrt{10}$, and $\epsilon=0.01$, and solve the trust region subproblem in the same way as the trust region Newton-conjugate gradient method (Algorithm 4.1 in Curtis2021TrustN ). For the ILMNegCur algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $L_{2}=50$, $\epsilon=0.01$, and solve the minimum eigenvalue subproblem by the Lanczos method, and the linear system subproblem by the conjugate gradient method. For the the IMINIMAX-TR algorithm, we set $N_{t}=30$, $\eta_{y}=0.01$, $L_{2}=50$, $\epsilon=0.01$, and solve the trust-region subproblem the same as in the IGRTR algorithm. For the GDA algorithm, we set the step size of both $x$ and $y$ to 0.01. In addition, for the IMCN algorithm, we solve the subproblem about ${y}$ the same as the IGRTR algorithm and the ILMNegCur algorithm. In addition, the cubic subproblem is solved by Algorithm 4 in Luo2022FindingSS , where $L=2$, $M=50$, $\sigma=0.001$, and $\mathcal{K}(\epsilon,\delta^{{}^{\prime}})=10$.
Figure 3 shows the numerical performance of the five tested algorithms when solving problem (4.3), where in sub-figure (a), the horizontal axis represents the running time and the vertical axis represents the test accuracy; while in sub-figure (b), the horizontal axis represents the number of iterations $t$ and the vertical axis represents the function value $P(x)$. It can be seen that the number of iterations and convergence time of all second-order algorithms are significantly less than that of the GDA algorithm, while IGRTR performs slightly better than other second-order algorithms. These results show the advantages of the two proposed algorithm.
@@ 5 Conclusions
In this paper, we propose a gradient norm regularized trust-region (GRTR) algorithm and a Levenberg-Marquardt algorithm with a gradient norm regularization coefficient and use the negative curvature direction to correct the iteration direction (LMNegCur). Both the GRTR algorithm and the LMNegCur algorithm can achieve an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$. The inexact variants of both methods require calling $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$ times of Hessian-vector products and $\tilde{\mathcal{O}}(\ell^{2}\rho^{0.5}\mu^{-2}\epsilon^{-1.5})$ times of gradient ascent steps to find $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary points with high probability. It is worthwhile to further study the application of random subsampling to further improve the performance and complexity of the proposed algorithm for large-scale nonconvex minimax optimization problems.